How many moles of an ideal gas at a temperature of -13 degree and a pressure of 2 atm will occupy a volume of 21.32 dm3 use R=8.31 Jmol-1 K-1

1
Expert's answer
2021-03-08T06:04:46-0500

To determine the number of moles, the use of the ideal gas low is required:

"pV=nRT"

where "p" is a pressure (Pa), "V" is a volume ("m^3"), "n" is a number of moles, "R" is a universal gas constant ("J/K\times mol") and "T" is a temperature (K).

From this law, the number of moles computes in the following way:

"n=\frac{RT}{pV}"

Taking into account the correct units, "n" equals to:

"n=\frac{8.31 J/K\times mol\cdot(273-13)K}{(2\cdot101325)Pa \cdot (21.32\cdot 0.001)m^3}=\frac{2160.6}{4320.5}=0.5 mol"